The fundamental frequency of an open pipe is 450 Hz and that of a closed pipe is 350 Hz. The two pipes are joined together to form a longer pipe. Find the fundamental frequency of this new pipe . Take velocity of sound as 330m/s

Grade:10

1 Answers

venkat
6 years ago
The length of closed pipe is given by    l_{1}=v/4n_{1} 
Similarly the length of open pipe  is given by l_{2}=v/2n_{2}
Now effective length of the closed pipe is given by l=l​1+l2
The fundamental frequency of the new closed pipe is given as 
n`=v/4(l_{1}+l_{2})
Please approve my answer if it helped you.